A steady current of 0.5A flows through a wire.

a. How much charge passes through it in one hour?

b. What is the number of electrons that pass through the wire during this time?


1
Expert's answer
2022-03-23T13:52:52-0400

(a)

"q=It\\
=\rm 0.5\: A\times 3600\: s=1800\: C"

(b)

"N=\frac{1800\:\rm C}{1.6\times 10^{-19}\:\rm C}=1.13\times10^{22}"